# Designing metasurface optical interfaces for solid-state qubits using many-body adjoint shape optimization

### v1

@@ I Introduction
Solid-state quantum defects — exemplified by the nitrogen-vacancy (NV) center in diamond—have emerged at the forefront of quantum technologies in recent years, seeing applications in core research areas of quantum computing Awschalom et al. (2018); Abobeih et al. (2022), communication Bradley et al. (2019); Pompili et al. (2021); Bersin et al. (2024), and sensing Zhou et al. (2020); Bürgler et al. (2023); Li et al. (2022). Measurements are performed optically, with the defect’s spin state being encoded in its single-photon photoluminescence. As a result, efficient photon capture is essential, with readout fidelity, sensitivity, and operational speed being limited by optical losses. One fundamental source of optical loss is diamond’s high refractive index at optical wavelengths (around 2.4), which causes the majority of emitted photons to be trapped at the flat air-diamond interface due to total internal reflection. Past works have addressed this challenge by embedding NV centers in nanophotonic structures such as waveguides, nanopillars, and photonic crystal cavitiesMouradian et al. (2015); Neu et al. (2014); Barclay et al. (2009); Chakravarthi et al. (2023); Zheng et al. (2023); such structures in the near-field overcome total internal reflection and increase the photon emission rate through Purcell enhancement. However, near-field structures also situate NV centers near diamond surfaces, which conversely degrade their crucial spin and optical features Ofori-Okai et al. (2012); Rosskopf et al. (2014); Chou et al. (2023).
For applications such as quantum memories in which maximizing the spin coherence time is paramount, it is necessary to use NV centers embedded deeper in the bulk diamond and to only modify the far-field propagation of emitted photons. The prototypical structure used in this scenario is the solid immersion lens (SIL), which consists of a hemisphere centered around a defect, such that any emitted photon will be incident normal to the air-diamond interface and not experience total internal reflection. These hemispheres are typically etched into the diamond Castelletto et al. (2011) but can also be created with additive fabrication methods Cheng et al. (2023). However, SILs require a difficult and time-consuming volumetric fabrication process, and they do not adjust the propagation direction of emitted light, so high numerical aperture (NA) free-space microscope objectives are still required to collimate the diverging beam.
In contrast, flat optics etched on the diamond surface can both transmit and redirect incident photons Li et al. (2015); Huang et al. (2019). Flat optics are easier and cheaper to fabricate in diamond than volumetric structures like SILs, making them an attractive option for scaled-up quantum computation or quantum communication systems. In addition, by simultaneously collimating the emitted photons, a flat optic can circumvent the need for expensive, bulky, free-space optics and ultimately allow the photons to be collected by an optical fiber in a compact, robust, inexpensive device.
@@ II.1 Background
Figure 1: (a) To-scale render of optimized collimating metasurface. (b) Top-down view of nanopillar array. Inset: Zoom-in on one section, showing the parameterization of each nanopillar. Since the center points of each ellipse are free parameters, the structure is not constrained to a fixed grid like most conventional metasurfaces, but constraint functions enforce a minimum distance between adjacent elements for fabrication considerations. (c) Diagram of simulations run during each step of adjoint optimization. Two forward and adjoint simulation pairs correspond to two different dipole orientations (orange arrows) that are mutually orthogonal to each other and to the [111] NV axis. This set of simulations determines a figure of merit and its gradient, which is fed into a gradient-descent optimization algorithm
@@
We utilize an inverse design method based on the adjoint method, which has been utilized extensively in the photonics field in recent years for applications varying in complexity from simple components to arbitrary equation solvers Molesky et al. (2018); Estakhri et al. (2019); Cordaro et al. (2023). With just two physics simulations — a typical “forward” simulation of optical propagation and an “adjoint” simulation that approximately represents the desired outcome sent backwards — an arbitrary figure of merit and its gradient can be calculated with respect to any number of design parameters. This can then be wrapped into a gradient-descent optimization algorithm to arrive at an optimized structure.
@@
Adjoint-based optimizations are categorized based on their parameterization method as either topology optimization or shape optimization. Most adjoint-based methods in recent years have made use of topology optimization, which treats the relative permittivity $\epsilon$ at each point in the optimization region as a free continuous parameter in order to explore a maximally broad parameter space. Typically, binarization and smoothing filters are applied gradually over the course of the optimization to ensure a realistic final structure made of discrete materials (e.g., using only two values of $\epsilon$) with smooth features Bendsøe and Sigmund (1999). Previous authors have used topology optimization to design optical metasurfaces based on patterned thin films Chung and Miller (2020); Lin et al. (2019); Wang et al. (2019) and nanophotonic structures using diamond membranes Dory et al. (2019). Topology optimization has also been utilized to design near-field antenna-like structures to extract photons from shallow NV centers Chakravarthi et al. (2020); Wambold et al. (2020).
@@
In contrast, shape optimization considers a fixed number of objects, whose boundaries are parameterized according to their size, position, orientation, etc,. Shape optimizations have been performed on individual objects to design structures such as Y-splitters Lalau-Keraly et al. (2013) and to generate unit cell libraries for larger metasurfaces Whiting et al. (2020). Other works have performed shape optimizations on periodically ordered arrays of objects Zhan et al. (2018); Mansouree et al. (2021) and for disordered photonic crystals Granchi et al. (2023).
@@ II.2 Implementation
The fabrication of diamond nanopillar arrays is well established. They are used, for example, in quantum sensing applications in which NV centers are embedded directly in the nanopillars Zou et al. (2008); Hausmann et al. (2011); Losero et al. (2023). Previous work has also shown how diamond nanopillar arrays can function as a metalens to collect light from deep NV centers Huang et al. (2019), motivating the present work. The device efficiency and NA of the metalens in Ref. Huang et al. (2019) was limited by the classical design strategy, which treated each pillar as a localized phase-shifting element.
The explicit formulation of the metasurface as an array of nanopillars ensures a fabricable design at each iteration of the optimization process. Fabrication constraints are enforced through lower bounds on the axial lengths to ensure a minimum pillar size as well as differentiable inequality constraint functions between nearest-neighbor pillars to ensure a minimum separation, $d$:
@@
Since there is no analytical expression for the distance between ellipses, the elliptical nanopillars are instead conservatively approximated as circles with $r_{i}=\max(a_{i},b_{i})$ for the constraint calculations. To perform the optimizations, we use these constraint functions along with adjoint-calculated gradients in an open-source implementation of the Method of Moving Asymptotes algorithm Svanberg (2002); Johnson (2007).
@@ II.3 Figure of Merit
We construct the figure of merit for adjoint optimization to maximize the percentage of emitted photons that escape the diamond surface into a desired collection angle. We consider the full NV emission spectrum, which at room temperature is dominated by a wide phonon sideband spanning $\approx$650-750 nm.
At a given collection plane above the NV surface, the propagation direction of the forward fields can be calculated by taking a spatial Fourier transform of the electric and magnetic fields. The angular acceptance constraint is equivalent to a circular region in $k$-space given by:
$R=\{k_{x},k_{y}:\sqrt{k_{x}^{2}+k_{y}^{2}}/k_{0}\leq\mathrm{NA_{target}}\}$ (2)
where $k_{x}$ and $k_{y}$ are the $x$ and $y$ wavevector components, $k_{0}=\frac{2\pi}{\lambda_{0}}$, and $\mathrm{NA_{target}}$ is the numerical aperture corresponding to the desired collection angle. For a given dipole orientation, $i$, the directional transmission into $\mathrm{NA_{target}}$ is given by integrating the time-averaged Poynting vector over this region in k-space:
@@
where $P_{\mathrm{dipole}}$ is the total power injected into the simulation and $\mathcal{F}[\mathbf{E}_{i}]$ and $\mathcal{F}[\mathbf{H}_{i}^{*}]$ are the spatial Fourier-transformed electric and magnetic fields on a plane above the diamond surface. $R$ in this case is the region defined in Eqn. 2 but could be any arbitrary region in k-space. The total figure of merit (FOM) is subsequently constructed as follows:
@@
We consider a diamond with a top surface aligned with a (100) crystal plane, which is typical of synthetic diamond plates used in NV-center experiments. The NV center itself is oriented along a $\langle 111\rangle$ axis, and its emission pattern results from an incoherent sum of two optical dipoles orthogonal to each other and to the NV axis, as illustrated in Fig. 1c. Less commonly, diamond can be grown Tallaire et al. (2014) or cleaved Parks et al. (2018) to create a (111) top surface. The (111)-oriented diamond hosts NV centers whose optical dipoles are oriented parallel to the surface and subsequently achieves higher efficiency. The optimization process can be easily performed under the assumption of (111) diamond; see Supplement 1 for a comparison of the final results.
The source for the adjoint simulation is derived from the figure of merit, and the gradients are calculated by integrating the forward and adjoint fields over the nanopillar surfaces according Eqns. (S1-S5) in the Supplement; see also Ref. Miller (2013). The simulations required to perform this optimization are depicted in Fig. 1c.
In order to simultaneously optimize over a broad frequency range, we perform simulations in the time domain using Lumerical FDTD Lumerical Inc. . The optimization procedure is implemented using the LumOpt package Lalau-Keraly et al. (2013) as a base for additional custom Python code, which implements the geometry construction, figure of merit and gradient calculations, and optimization algorithm wrapper (See Code 1, Ref. Klein (2023)).
@@ III.1 Performance
Figure 2: (a) Performance metric of final structure. Total photon transmission efficiency is calculated as a function of the numerical aperture of an ideal collection optic placed above the metasurface. The dashed line indicates the target used in the optimization. The right y-axis normalizes the transmission by the solid angle of the directional NV emission incident on the surface. (b) Performance is plotted as a function of wavelength for directional transmission efficiency (within $\mathrm{NA_{col}}$ = 0.2) and maximum transmission (within $\mathrm{NA_{col}}$ = 1). The gray line shows a normalized spectrum of a NV center measured at room temperature in Huang et al. (2019) for weighting the spectral response to calculate the figure of merit.
@@ III.2 Effect of Initial Conditions
The second pre-selected parameter was the initial spacing of nanopillars, which determines the total number of nanopillars present in the optimization. Although the nanopillars are allowed to move freely during the optimization process, the initial condition for the structure discussed earlier was a regular grid of nanopillars with a spacing of 300 nm. Over a 5 µm x 5 µm area, this corresponds to a total of 256 nanopillars, or 1280 optimization parameters. No nanopillars were added or removed over the course of the optimization in order to maintain a fixed number of optimization parameters. In principle, the optimization procedure could be adapted to allow for the addition or removal of nanopillars between optimization phases.
In order to determine the effect of nanopillar height and spacing, we performed several additional optimizations varying these pre-selected values. The results are quantified in Fig. 4, and the final geometries from each optimization are shown in Figs. S3-S4 in the supplementary information. In order to account for variations in convergence, we performed five optimizations for each initial condition in order to quantify the range of potential outcomes. The five optimizations used identical starting grids but varied the initial axial lengths of the nanopillars. In one optimization, each nanopillar was initialized with a 100-nm-diameter circular cross-section; for the other four, the axial lengths were initially set to random values between 25 nm and 125 nm. We did not observe any trend as to whether the uniform or randomized starting axial lengths led to better-performing final structures.
@@ III.3 Tolerance to Fabrication Errors
We simulated the effects of potential fabrication errors and plotted the results in Fig. 5. The performance losses are expressed in units of decibels (dB) compared to the performance as designed. Figures 5a and 5b show the performance loss due to lateral and vertical misalignment, respectively. The loss is calculated by performing many separate FDTD simulations with the emitter placed at varying positions. For lateral misalignments, the 3 dB loss contour occurs for displacements around 470 nm. The structure shows similar sensitivity to displacements in the emitter depth. With the design targeted to a depth of 1 µm beneath the metasurface, we predict a loss within 3 dB with the emitter instead placed as shallow as 0.35 µm or as deep as 1.35 µm. Lithographic fabrication methods can be used to align the metasurface over a chosen NV center well within these fabrication tolerances — previous work has successfully aligned solid immersion lenses onto pre-selected NV centers of a particular depth with lateral accuracy less than 100 nm and higher accuracy is possible Jamali et al. (2014). If desired, the alignment sensitivity could be further decreased by explicitly incorporating it into the optimization process by co-optimizing the performance from multiple emitter locations.

### v2

@@ I Introduction
Solid-state quantum defects — exemplified by the nitrogen-vacancy (NV) center in diamond—have emerged at the forefront of quantum technologies in recent years, seeing applications in core research areas of quantum computing [1, 2], communication [3, 4, 5], and sensing [6, 7, 8]. Measurements are performed optically, with the defect’s spin state being encoded in its single-photon photoluminescence. As a result, efficient photon capture is essential, with readout fidelity, sensitivity, and operational speed being limited by optical losses. One fundamental source of optical loss is diamond’s high refractive index at optical wavelengths (around 2.4), which causes the majority of emitted photons to be trapped at the flat air-diamond interface due to total internal reflection. Past works have addressed this challenge by embedding NV centers in nanophotonic structures such as waveguides, nanopillars, and photonic crystal cavities[9, 10, 11, 12, 13]; such structures in the near-field overcome total internal reflection and increase the photon emission rate through Purcell enhancement. However, near-field structures also situate NV centers near diamond surfaces, which conversely degrade their crucial spin and optical features [14, 15, 16].
For applications such as quantum memories in which maximizing the spin coherence time is paramount, it is necessary to use NV centers embedded deeper in the bulk diamond and to only modify the far-field propagation of emitted photons. The prototypical structure used in this scenario is the solid immersion lens (SIL), which consists of a hemisphere centered around a defect, such that any emitted photon will be incident normal to the air-diamond interface and not experience total internal reflection. These hemispheres are typically etched into the diamond [17] but can also be created with additive fabrication methods [18]. However, SILs require a difficult and time-consuming volumetric fabrication process, and they do not adjust the propagation direction of emitted light, so high numerical aperture (NA) free-space microscope objectives are still required to collimate the diverging beam.
In contrast, flat optics etched on the diamond surface can both transmit and redirect incident photons [19, 20]. Flat optics are easier and cheaper to fabricate in diamond than volumetric structures like SILs, making them an attractive option for scaled-up quantum computation or quantum communication systems. In addition, by simultaneously collimating the emitted photons, a flat optic can circumvent the need for expensive, bulky, free-space optics and ultimately allow the photons to be collected by an optical fiber in a compact, robust, inexpensive device.
@@ II.1 Background
Figure 1: (a) To-scale render of optimized collimating metasurface. (b) Top-down view of nanopillar array. Inset: Zoom-in on one section, showing the parameterization of each nanopillar. Since the center points of each ellipse are free parameters, the structure is not constrained to a fixed grid like most conventional metasurfaces, but constraint functions enforce a minimum distance between adjacent elements for fabrication considerations. (c) Diagram of simulations run during each step of adjoint optimization. Two forward and adjoint simulation pairs correspond to two different dipole orientations (orange arrows) that are mutually orthogonal to each other and to the [111] NV axis. This set of simulations determines a figure of merit and its gradient, which is fed into a gradient-descent optimization algorithm.
@@
We utilize an inverse design method based on the adjoint method, which has been utilized extensively in the photonics field in recent years for applications varying in complexity from simple components to arbitrary equation solvers [21, 22, 23]. With just two physics simulations — a typical ”forward” simulation of optical propagation and an ”adjoint” simulation that approximately represents the desired outcome sent backwards — an arbitrary figure of merit and its gradient can be calculated with respect to any number of design parameters. This can then be wrapped into a gradient-descent optimization algorithm to arrive at an optimized structure.
@@
Adjoint-based optimizations are categorized based on their parameterization method as either topology optimization or shape optimization. Most adjoint-based methods in recent years have made use of topology optimization, which treats the relative permittivity $\epsilon$ at each point in the optimization region as a free continuous parameter in order to explore a maximally broad parameter space. Typically, binarization and smoothing filters are applied gradually over the course of the optimization to ensure a realistic final structure made of discrete materials (e.g., using only two values of $\epsilon$) with smooth features [24]. Previous authors have used topology optimization to design optical metasurfaces based on patterned thin films [25, 26, 27] and nanophotonic structures using diamond membranes [28]. Topology optimization has also been utilized to design near-field antenna-like structures to extract photons from shallow NV centers [29, 30].
@@
In contrast, shape optimization considers a fixed number of objects, whose boundaries are parameterized according to their size, position, orientation, etc,. Shape optimizations have been performed on individual objects to design structures such as Y-splitters [31] and to generate unit cell libraries for larger metasurfaces [32]. Other works have performed shape optimizations on periodically ordered arrays of objects [33, 34] and for disordered photonic crystals [35, 36]. Additionally, we became aware of similar works performing inverse design on metasurfaces using freely-moving pillars while this manuscript was under review [37, 38].
@@ II.2 Implementation
The fabrication of diamond nanopillar arrays is well established. They are used, for example, in quantum sensing applications in which NV centers are embedded directly in the nanopillars [39, 40, 41]. Previous work has also shown how diamond nanopillar arrays can function as a metalens to collect light from deep NV centers [20], motivating the present work. The device efficiency and NA of the metalens in Ref. [20] was limited by the classical design strategy, which treated each pillar as a localized phase-shifting element.
The explicit formulation of the metasurface as an array of nanopillars ensures a fabricable design at each iteration of the optimization process. Fabrication constraints are enforced through lower bounds on the axial lengths to ensure a minimum pillar size as well as differentiable inequality constraint functions between nearby pillar pairs $(P_{i},P_{j})$ to ensure a minimum separation, $d$:
@@
Since there is no analytical expression for the distance between ellipses, the elliptical nanopillars are instead conservatively approximated as circles with $r_{i}=\max(a_{i},b_{i})$ for the constraint calculations. These constraint functions are applied between adjacent pillars based on their initial configuration on a rectangular grid, including vertically, horizontally, and diagonally adjacent pairs. While pillars are allowed to move freely over the course of the optimization, constraints based only on their initial positions proved sufficient for enforcing a minimum gap size in the optimizations shown in this work. To perform the optimizations, we use these constraint functions along with adjoint-calculated gradients in an open-source implementation of the Method of Moving Asymptotes algorithm [42, 43].
@@ II.3 Figure of Merit
We construct the figure of merit for adjoint optimization to maximize the percentage of emitted photons that escape the diamond surface into a desired collection angle. We consider the full NV emission spectrum, which at room temperature is dominated by a wide phonon sideband spanning $\approx$650-750 nm. At a collection plane above the NV surface, the simulated electric and magnetic fields are expressed as a function of propagation directions by taking a spatial Fourier transform. These fields are filtered according to the function $u(k_{x},k_{y})$ to include only the fields within the acceptance angle of a large, multi-mode fiber with a numerical aperture of $\mathrm{NA_{target}}$ placed above the metasurface:
$u(k_{x},k_{y})=\begin{cases}1&\sqrt{k_{x}^{2}+k_{y}^{2}}/k_{0}\leq\mathrm{NA_{target}}\\ 0&\mathrm{else}\end{cases}$ (2)
where $k_{x}$ and $k_{y}$ are the $x$ and $y$ wavevector components and $k_{0}=\frac{2\pi}{\lambda_{0}}$. For a given dipole orientation i, the filtered electric fields are described by applying this filter on the Fourier-transformed fields and then taking an inverse Fourier transform to return to real space:
$\mathbf{E}_{\mathrm{filt}}^{i}(\mathbf{r})=\mathcal{F}^{-1}\Big{[}{u(k_{x},k_{y})\mathcal{F}[\mathbf{E}^{i}(\mathbf{r})]}\Big{]}$ (3)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ describe the spatial Fourier transform and its inverse, and the filtered magnetic fields $\mathbf{H}_{\mathrm{filt}}^{i}(\mathbf{r})$ are calculated analogously. These real-space filtered fields are needed to calculate the source for the adjoint simulations as described in Supplement 1 and are also integrated over the collection plane to obtain the directional transmission into $\mathrm{NA_{target}}$:
@@
where $P_{\mathrm{dipole}}$ is the total power injected into the simulation. The total figure of merit (FOM) is subsequently constructed as follows:
@@
We consider a diamond with a top surface aligned with a (100) crystal plane, which is typical of synthetic diamond plates used in NV-center experiments. The NV center itself is oriented along a $\langle 111\rangle$ axis, and its emission pattern results from an incoherent sum of two optical dipoles orthogonal to each other and to the NV axis, as illustrated in Fig. 1c. Less commonly, diamond can be grown [44] or cleaved [45] to create a (111) top surface. The (111)-oriented diamond hosts NV centers whose optical dipoles are oriented parallel to the surface and subsequently achieves higher efficiency. The optimization process can be easily performed under the assumption of (111) diamond; see Supplement 1 for a comparison of the final results.
The source for the adjoint simulation is derived from the figure of merit, and the gradients are calculated by integrating the forward and adjoint fields over the nanopillar surfaces according Eqns. (S1-S5) in the Supplement; see also Ref. [46]. The simulations required to perform this optimization are depicted in Fig. 1c.
In order to simultaneously optimize over a broad frequency range, we perform simulations in the time domain using Lumerical FDTD [47]. The optimization procedure is implemented using the LumOpt package [31] as a base for additional custom Python code, which implements the geometry construction, figure of merit and gradient calculations, and optimization algorithm wrapper (See Code 1, Ref. [48]).
@@ III.1 Performance
Figure 2: (a) Performance metric of final structure. Total photon transmission efficiency is calculated as a function of the numerical aperture of an ideal collection optic placed above the metasurface. The dashed line indicates the target used in the optimization. The right y-axis normalizes the transmission by the solid angle of the directional NV emission incident on the surface. (b) Performance is plotted as a function of wavelength for directional transmission efficiency (within $\mathrm{NA_{col}}$ = 0.2) and maximum transmission (within $\mathrm{NA_{col}}$ = 1). The gray line shows a normalized spectrum of a NV center measured at room temperature in [20] for weighting the spectral response to calculate the figure of merit.
@@
Additionally, we demonstrate that our choice in allowing the position of pillars to be free parameters during the optimization results in a final design with a higher performance than one constrained to pillars fixed on a unit-cell grid. We do this by performing an additional shape optimization in which the pillar positions are no longer free parameters, and obtain a directional transmission efficiency of 2.56%, which is a 29% decrease from the case in which pillars are allowed to move. Both structures and their performance are plotted together in the supplementary information in Fig. S1.
@@ III.2 Effect of Initial Conditions
The second pre-selected parameter was the initial spacing of nanopillars, which determines the total number of nanopillars present in the optimization. Although the nanopillars are allowed to move freely during the optimization process, the initial condition for the structure discussed earlier was a regular grid of nanopillars with a spacing of 300 nm. Over a 5 µm x 5 µm area, this corresponds to a total of 256 nanopillars, or 1280 optimization parameters. No nanopillars were added or removed over the course of the optimization in order to maintain a fixed number of optimization parameters.
In principle, the optimization procedure could be expanded upon to allow for the addition or removal of nanopillars between optimization phases by utilizing the gradients and constraint functions that are already calculated at each iteration. For example, pillars could be removed if their derivative with respect to both axial lengths are negative but the lengths themselves are already at their lower bound determined by the minimum feature size. Pillar pairs could be combined if their derivatives indicate they should be closer together but they are limited by the constraint functions described in Eqn. 1. And new pillars could be added if the gradient fields, described as the dot product between forward and adjoint electric fields (see Eqn. S1), are large enough to indicate a pillar should exist in an empty region.
In order to determine the effect of nanopillar height and spacing, we performed several additional optimizations varying these pre-selected values. The results are quantified in Fig. 4, and the final geometries from each optimization are shown in Figs. S4-S5 in the supplementary information. In order to account for variations in convergence, we performed five optimizations for each initial condition in order to quantify the range of potential outcomes. The five optimizations used identical starting grids but varied the initial axial lengths of the nanopillars. In one optimization, each nanopillar was initialized with a 100-nm-diameter circular cross-section; for the other four, the axial lengths were initially set to random values between 25 nm and 125 nm. We did not observe any trend as to whether the uniform or randomized starting axial lengths led to better-performing final structures.
@@
Overall, we find that the choice of initial pillar pitch does not strongly impact the final device performance as long as the total number of parameters is not too small (as with the 400 nm batch) and the dynamic range of values for each parameter is not limited by constraints (as with the 200 nm batch). Additionally, the choice of a 300 nm pitch and 750 nm height we chose based on this analysis is very similar to the choice of a 300 nm pitch and 1 µm height chosen to achieve a full $2\pi$ phase response to design the phase-gradient metasurface in Ref. [20], with the 750 nm height performing only marginally better in the best-case scenario. We expect that in general, choosing fixed parameters similar to those used for a conventionally-designed device should lead to good results.
@@ III.3 Tolerance to Fabrication Errors
We simulated the effects of potential fabrication errors and plotted the results in Fig. 5. The performance losses are expressed in units of decibels (dB) compared to the performance as designed. Figures 5a and 5b show the performance loss due to lateral and vertical misalignment, respectively. The loss is calculated by performing many separate FDTD simulations with the emitter placed at varying positions. For lateral misalignments, the 3 dB loss contour occurs for displacements around 470 nm. The structure shows similar sensitivity to displacements in the emitter depth. With the design targeted to a depth of 1 µm beneath the metasurface, we predict a loss within 3 dB with the emitter instead placed as shallow as 0.35 µm or as deep as 1.35 µm. Lithographic fabrication methods can be used to align the metasurface over a chosen NV center well within these fabrication tolerances — previous work has successfully aligned solid immersion lenses onto pre-selected NV centers of a particular depth with lateral accuracy less than 100 nm and higher accuracy is possible [49]. If desired, the alignment sensitivity could be further decreased by explicitly incorporating it into the optimization process by co-optimizing the performance from multiple emitter locations.